# Non-commutative effective field theory of the lowest Landau level superfluid

### v2

@@ Abstract
A 2+1D superfluid in a rapidly rotating trap forms an array of vortices, with collective excitations called Tkachenko modes. Reference 1 argued from an effective field theory viewpoint that these excitations are described by a field theory living on a non-commutative space. We elucidate the microscopic origin of these non-commutative fields, and present a novel derivation of the effective field theory for this superfluid using a lowest Landau level projected coherent state path integral approach. Besides conceptual clarity, this approach makes quantitative predictions about the long-wavelength, low-energy behavior in terms of the microscopic parameters of the short-range interacting lowest Landau level superfluid — relevant to trapped Bose-Einstein condensate experiments.
@@ I Introduction
Following the discovery of the fractional quantum Hall effect [2], it has been appreciated that two-dimensional quantum gases in strong magnetic fields produce exotic phases of matter originating from strong interactions engendered by quenched kinetic energy. Since then, a diverse set of approaches describing these states have emerged, including trial wavefunctions [3], Haldane pseudopotentials [4], flux attachment [5, 6], Chern-Simons theory [7] for the gapped fractional quantum Hall states, and Halperin-Lee-Read (HLR) theory [8, 9] for the metallic phase of the half-filled Landau level. Field theoretic descriptions of these states such as HLR and Chern-Simons theory are notably not manifestly restricted to the lowest Landau level (LLL), even though the higher Landau levels are gapped out and, for the subset of these states that we are interested in, the relevant physics happens completely within the LLL as confirmed by numerical simulations [10]. A natural question to ask is whether we can write down field theories in such a way that the LLL projection is manifest.
@@
Non-commutative fields have been claimed to be useful to describe physics manifestly in the LLL since the late 1990s [11, 12, 13, 14, 15, 16, 17] (for a review see Ref. [18]). Recent advances have exploited this technology to describe the $\nu=1$ bosonic composite Fermi liquid state [19, 20, 21], which uses an exact Pasquier-Haldane map [22] to develop the non-commutative field theory. It has also been used to describe traditional Fermi liquids, where phase space is treated as a non-commutative space [23] . Another work [1] studied the LLL superfluid using boson-vortex duality [24] and argues that non-commutative field theories are the right language that describes the fluctuations above the mean-field vortex configuration. Using the non-commutative field theory, one can derive the correct scaling behavior of the dispersion relation and the decay rate of the low-energy excitations of this system, called Tkachenko modes [25, 26, 27, 28, 29, 30, 31]. The low energy theory is described by a quantum Lifshitz model (a similar theory has been used to describe vortex lattices in layered superconductors [32]). Although the mean-field vortex lattice state breaks both translation symmetry and $U(1)$ charge conservation symmetry spontaneously, this state only has one branch of low energy excitations, which disperses quadratically as $\omega\sim k^{2}$ [33].
@@
This article solves the problem of how to arrive at a manifestly LLL-projected effective field theory (EFT) for the superfluid phase of bosons in the LLL, which can be experimentally realized in systems of Bose-Einstein condensates (BECs) in a rapidly rotating harmonic trap [34, 35, 36]. The problem of a rapidly rotating superfluid with a fine-tuned harmonic trap can be mapped at low energies to a LLL superfluid with an effective cyclotron frequency twice the trap rotation frequency. We refer to earlier works for readers interested in the explicit mapping [37]. In this work, we abstract away to the problem of bosons restricted to the LLL interacting via a contact repulsive interaction and derive an effective field theory for the LLL superfluid phase.
@@
Starting from a generic LLL-projected Hamiltonian for bosons describing a superfluid, we derive the non-commutative field theory of Ref. 1 using physically motivated assumptions. We derive this effective description by developing a manifestly LLL-projected coherent state path integral. This approach both allows us to justify the non-commutative effective field theory for the LLL superfluid, and helps us connect these abstract, effective degrees of freedom to the microscopic contituents. In particular, for systems of BECs in rotating traps, which are composed of weakly interacting bosons, one can calculate the coefficients of the non-commutative Lagrangian in terms of the microscopic parameters and get quantitative predictions for the dispersion relation and decay rate of Tkachenko modes. We calculate the dispersion relation in this paper and leave the calculation of decay rate to future work.
@@
The outline of the paper goes as follows: in Section II, we introduce the single-particle physics of the lowest Landau level and develop the machinery which we later use to define coherent states in Section III. In the process, we introduce the mathematics of non-commutative (NC) functions and refer the reader to Appendix B for a more thorough discussion. In Section III, we develop the coherent state path integral restricted to the LLL that captures the effective degrees of freedom of the LLL superfluid, and in Section IV effectively derive the non-commutative field theoretic action described in Ref. 1, using some symmetry arguments and dimensional analysis to arrive at the form of the most relevant interactions.
@@ II Single particle physics in the LLL
Due to the nontrivial commutation relations of the guiding center operators, these magnetic translation operators satisfy a closed algebra known as the Girvin-Macdonald-Platzman (GMP) algebra [38]. This can be derived using the Baker-Campbell-Hausdorff formula and is given by
@@
Simultaneously diagonalizing these operators allows us to find magnetic lattice translation invariant wavefunctions where the zeros of the holomorphic function $f(z)$ are arranged in a Bravais lattice defined by the translation vectors $\bm{a}_{i}=l_{B}^{2}(\wedge\bm{k}_{i})$. We can also interpret the zeros of the wavefunctions as vortices, since the wavefunction winds by $2\pi$ as one traverses around it. So, we will call such wavefunctions vortex lattices. The specific form of the holomorphic functions is irrelevant for the discussion – they can be constructed using elliptic theta functions or Weierstrass sigma functions [39]. One useful and intuitive construction of these functions is using the strip-like wavefunctions. For the case of square lattice $\bm{a}_{1}=(\sqrt{2\pi}l_{B},0),\bm{a}_{2}=(0,\sqrt{2\pi}l_{B})$, the wavefunction $|\Phi_{0}\rangle$} with eigenvalue one for both $\hat{T}_{1}$ and $\hat{T}_{2}$ can be written as
@@
due to linearlity (the $\approx$ denotes coarse-graining and ignoring higher-order terms in $\left|\bm{k}\right|^{2}l_{B}^{2}$). This motivates a mathematical map between NC functions and regular functions, known as the Weyl map [18]. The Weyl map is a linear, invertible map given by
@@ III Many body bosonic coherent states in the LLL
Note that the projection to the LLL comes for free as the coherent states are LLL-restricted, and that the NC field $\hat{f}$ describes a boson. The last step is paramount. The time derivative and chemical potential terms match exactly when we rewrite it in terms of the trace over non-commutative space, but the quartic term cannot be exactly mapped in that way. From the long-wavelength behavior of the wavefunction products (Appendix A and discussions earlier), we know that it is approximated by the leading order term written above, where $g^{\ast}(\approx 1.160g$ for the triangular vortex lattice [40], details in Appendix C) is a renormalized interaction vertex, which can be computed by numerically estimating the interaction energy of the reference state $\sim\int d^{2}\bm{r}\left|\Phi_{0}(\bm{r})\right|^{4}$. Since this is not an exact mapping, there are corrections that have to be quartic in the field operators because of the scaling with density, and they come in the form of higher derivative terms in the NC field theory language. These corrections depend on the exact vortex lattice reference state $|\Phi_{0}\rangle$} that we start with, and we will not try to calculate the coefficients from microscopics. Rather, we use symmetry and dimensional analysis to constrain the form of the correction terms.
@@
We find that there is no derivative four-field term that both respect the spatial symmetries and gives an energy cost to unitary $\hat{f}$. There is a similar issue if one tries to write down a local field theory in terms of the boson field operators for a dipole conserving Bose-Hubbard model [41] which captures the effects of dipole hopping. Another approach is to start from a non-local [^3] Non-local here just means that the terms in the Lagrangian are not simply a function of a field and its derivatives at the same spatial point. The theory is still physically local, it has exponential decay of interactions. hopping term in the Lagrangian
@@
where $g(\bm{a},\bm{b})=g\delta(\bm{a}\cdot\bm{b})\mathcal{O}(e^{-(\bm{a}^{2}+\bm{b}^{2})l_{B}^{-2}})$ is a decaying function reflecting the locality of the hopping process. This form of the hopping term is constrained by magnetic translation and rotation symmetry, which implies dipole moment $\int\rho\hat{r}_{i}$ and trace of quadrupole moment $\int\rho\hat{r}_{i}\hat{r}_{i}/2$ conservation respectively, as these are the generators of those symmetries [42]. Then, we use Hubbard-Stratonovich fields (Appendix B) to resolve the interaction terms into quadratic pieces. Because of the symmetries of the problem, we have to introduce two sets of Hubbard-Stratonovich fields, one for the trace of quadrupole moment and a vector field for the dipole moment. Using this, we can systematically form condensates of each of these fields. The mean-field phases naively correspond to
@@
    Crystalline phase: a phase with spontaneously broken rotation and translation symmetry, with a quadratically dispersing NGB. The naive expectation of two linearly dispersing NGB is subverted due to the symmetry allowed $\epsilon_{ij}d_{i}\partial_{\tau}d_{j}$ term, where $e^{id_{j}}$ is the two-component dipole/translation order parameter. This is similar to the NGB physics of a Heisenberg ferromagnet, and is a type-B NGB [33, 43].
@@
The naive mean-field theory predicts that there is a cascade of transitions between the trivial insulator and the LLL superfluid via spontaneous breaking of various higher moments of charge conservation symmetry. This naive treatment is wrong, and there is a simple reason why. Take the crystalline phase, for example. This phase can be thought of as a dipole condensate, that is, a phase where dipoles are proliferated starting from the parent trivial insulator. A necessary condition for this to be true is that the average charge density of this state is the same as the trivial insulator phase – which has zero density. But we know that the bosonic vacuum is the only state with exactly zero density, and there are no negative density states [^4] An identical issue arises around the zero filling Mott insulator lobe of the dipolar Bose-Hubbard model and is discussed in footnote 2 of [41].. Using an analogous argument for the hexatic phase (which is a condensate of quadrupoles), we conclude that there are no non-trivial hexatic and crystalline phases in this problem, and we generically expect a first-order transition between the trivial insulator phase and the LLL superfluid phase. Refs. [44, 45] study these phases from the dual perspective, i. e. in terms of melting of the vortex lattice. Under the identification of the vortex‐liquid with the trivial phase introduced above, their findings appear to contradict our results. Ref. [45] describes a series of transitions connecting the vortex solid (LLL superfluid) to the vortex liquid (trivial) phase, via the intermediate phases described above. This disagrees with our results as we argue that these intermediate phases do not exist, which becomes apparent when viewed from the original (non-dual) language. Ref. [44] speculates about a continuous direct transition between the vortex solid and vortex liquid phase, but we argue that such a transition will generically be first order. This is because from our perspective, the vortex liquid phase is the same as the bosonic vacuum, and one cannot have a continuous transition between states with different boson density. The disagreement is resolved when we note that the ideas in Refs. [44, 45] become relevant when we observe transitions between bosonic quantum Hall states and LLL superfluids, where the quantum Hall state is a vortex liquid phase at non-zero boson density, distinct from the bosonic vacuum state we consider here. We leave a microscopic understanding to the effective field theory describing this transition to future work.
@@ IV Non-commutative field theory of the LLL superfluid
With the form of the effective non-commutative Lagrangian for the interacting bosons in the LLL, we are ready to reproduce the effective NC field theory action for the LLL superfluid described in Ref. [1] by integrating out the amplitude mode. Like discussed earlier, we will use some symmetry and dimensional analysis arguments to arrive at the term that provides a dispersion for the Tkachenko mode.
@@
Focusing on the action, we note that the first term is invarant under time-reparametrization symmetry $\hat{u}(\tau)\mapsto\hat{u}(\alpha(\tau))$, where $\alpha$ is a differentiable map. This means that the first term is a geometric phase, since it only depends on the path we take and not how quickly we traverse the path. This is the non-commutative Berry Lagrangian term, discussed previously in a commutative field theoretic description of a LLL superfluid [46]. It encodes the Aharanov-Bohm phase picked up by the superfluid bosons due to the magnetic field. One way to see that is by noting that $\hat{u}$ is a unitary map in the LLL, which in the Heisenberg picture encodes an area preserving deformation (APD) of space [47], given by
@@
That is, we compute the energy cost $\int d^{2}\bm{r}gn_{0}^{2}\left|\Phi_{\hat{u}}(\bm{r})\right|^{4}$ numerically and match it with the effective field theory to get the correct coefficients. The details are shown in Appendix C.2. This quadratic low-energy action is valid as long as the system is short-range interacting, and at the thermodynamic and LLL limit. The Tkachenko field $\hat{\varphi}$ describes both the coarse-grained phase fluctuations of the LLL superfluid, and (its gradient describes) phonon fluctuations of the vortex lattice [1]. The Berry Lagrangian term is a total derivative term at quadratic order, and does not affect the dispersion relation of the Tkachenko mode, but this term is crucial for the transport physics of the LLL superfluid [45]. We have reproduced the quadratic EFT described in Ref. 1, along with its coefficients in terms of the microscopic parameters. The dispersion relation
@@
We end this section by noting that the quadratic dispersion of the Tkachenko mode implies that there is no long-range order of $e^{i\hat{\varphi}}$ in the 2+1D LLL superfluid, due to a generalized Hohenberg-Mermin-Wagner theorem [49, 50, 51, 42]. Instead, there is quasi-long range order and the field theory for the Tkachenko mode describes the spin-wave type excitations similar to the 1+1D superfluid, ignoring topological defects. Although the Tkachenko modes also describe the phonon motion of the vortex lattice (note that translation is spontaneously broken unlike $U(1)$ charge conservation), we believe it is inappropriate to call it a Nambu-Goldstone boson.

### v3

@@ Abstract
A 2+1D superfluid in a rapidly rotating trap forms an array of vortices, with collective excitations called Tkachenko modes. Reference Du2024a argued from an effective field theory viewpoint that these excitations are described by a field theory living on a non-commutative space. We elucidate the microscopic origin of these non-commutative fields, and present a novel derivation of the effective field theory for this superfluid using a lowest Landau level projected coherent state path integral approach. Besides conceptual clarity, this approach makes quantitative predictions about the long-wavelength, low-energy behavior in terms of the microscopic parameters of the short-range interacting lowest Landau level superfluid — relevant to trapped Bose-Einstein condensate experiments.
@@ I Introduction
Following the discovery of the fractional quantum Hall effect [ExptTsuiStormerGossard1982], it has been appreciated that two-dimensional quantum gases in strong magnetic fields produce exotic phases of matter originating from strong interactions engendered by quenched kinetic energy. Since then, a diverse set of approaches describing these states have emerged, including trial wavefunctions [Laughlin1983], Haldane pseudopotentials [Haldane1983], flux attachment [Wilczek1982, Jain1989], Chern-Simons theory [LopezFradkin1991] for the gapped fractional quantum Hall states, and Halperin-Lee-Read (HLR) theory [HLR1993, Son2015] for the metallic phase of the half-filled Landau level. Field theoretic descriptions of these states such as HLR and Chern-Simons theory are notably not manifestly restricted to the lowest Landau level (LLL), even though the higher Landau levels are gapped out and, for the subset of these states that we are interested in, the relevant physics happens completely within the LLL as confirmed by numerical simulations [BookPrange2012Hall]. A natural question to ask is whether we can write down field theories in such a way that the LLL projection is manifest.
@@
Non-commutative fields have been claimed to be useful to describe physics manifestly in the LLL since the late 1990s [Read1998, Polychronakos2001, SimeonHellerman2001, Susskind2001, Karabali2005, KarabaliNair2006, Pasquier2007] (for a review see Ref. [DouglasNekrasov2001]). Recent advances have exploited this technology to describe the $\nu=1$ bosonic composite Fermi liquid state [DongSenthil2020, DongSenthil2022, Goldman2022], which uses an exact Pasquier-Haldane map [PasquierHaldane1998] to develop the non-commutative field theory. It has also been used to describe traditional Fermi liquids, where phase space is treated as a non-commutative space [Padayasi2025] . Another work [Du2024a] studied the LLL superfluid using boson-vortex duality [DasguptaHalperin1981] and argues that non-commutative field theories are the right language that describes the fluctuations above the mean-field vortex configuration. Using the non-commutative field theory, one can derive the correct scaling behavior of the dispersion relation and the decay rate of the low-energy excitations of this system, called Tkachenko modes [Tkachenko1966a, Tkachenko1966b, Tkachenko1969, Sonin2014, Sonin1976, Volovik1979, Baym2003]. The low energy theory is described by a quantum Lifshitz model (a similar theory has been used to describe vortex lattices in layered superconductors [BalentsRadzihovsky1996]). Although the mean-field vortex lattice state breaks both translation symmetry and $U(1)$ charge conservation symmetry spontaneously, this state only has one branch of low energy excitations, which disperses quadratically as $\omega\sim k^{2}$ [WatanabeMurayama2013].
@@
This article solves the problem of how to arrive at a manifestly LLL-projected effective field theory (EFT) for the superfluid phase of bosons in the LLL, which can be experimentally realized in systems of Bose-Einstein condensates (BECs) in a rapidly rotating harmonic trap [BookCooperFQHBoson, ExptFletcherScience2021, ExptMukherjeeNature2022]. The problem of a rapidly rotating superfluid with a fine-tuned harmonic trap can be mapped at low energies to a LLL superfluid with an effective cyclotron frequency twice the trap rotation frequency. We refer to earlier works for readers interested in the explicit mapping [SinhaShlyapnikov2005]. In this work, we abstract away to the problem of bosons restricted to the LLL interacting via a contact repulsive interaction and derive an effective field theory for the LLL superfluid phase.
@@
Starting from a generic LLL-projected Hamiltonian for bosons describing a superfluid, we derive the non-commutative field theory of Ref. Du2024a using physically motivated assumptions. We derive this effective description by developing a manifestly LLL-projected coherent state path integral. This approach both allows us to justify the non-commutative effective field theory for the LLL superfluid, and helps us connect these abstract, effective degrees of freedom to the microscopic contituents. In particular, for systems of BECs in rotating traps, which are composed of weakly interacting bosons, one can calculate the coefficients of the non-commutative Lagrangian in terms of the microscopic parameters and get quantitative predictions for the dispersion relation and decay rate of Tkachenko modes. We calculate the dispersion relation in this paper and leave the calculation of decay rate to future work.
@@
The outline of the paper goes as follows: in Section II, we introduce the single-particle physics of the lowest Landau level and develop the machinery which we later use to define coherent states in Section III. In the process, we introduce the mathematics of non-commutative (NC) functions and refer the reader to Appendix B for a more thorough discussion. In Section III, we develop the coherent state path integral restricted to the LLL that captures the effective degrees of freedom of the LLL superfluid, and in Section IV effectively derive the non-commutative field theoretic action described in Ref. Du2024a, using some symmetry arguments and dimensional analysis to arrive at the form of the most relevant interactions.
@@ II Single particle physics in the LLL
Due to the nontrivial commutation relations of the guiding center operators, these magnetic translation operators satisfy a closed algebra known as the Girvin-Macdonald-Platzman (GMP) algebra [GMP1986]. This can be derived using the Baker-Campbell-Hausdorff formula and is given by
@@
Simultaneously diagonalizing these operators allows us to find magnetic lattice translation invariant wavefunctions where the zeros of the holomorphic function $f(z)$ are arranged in a Bravais lattice defined by the translation vectors $\bm{a}_{i}=l_{B}^{2}(\wedge\bm{k}_{i})$. We can also interpret the zeros of the wavefunctions as vortices, since the wavefunction winds by $2\pi$ as one traverses around it. So, we will call such wavefunctions vortex lattices. The specific form of the holomorphic functions is irrelevant for the discussion – they can be constructed using elliptic theta functions or Weierstrass sigma functions [HaldaneRezayi1985]. One useful and intuitive construction of these functions is using the strip-like wavefunctions. For the case of square lattice $\bm{a}_{1}=(\sqrt{2\pi}l_{B},0),\bm{a}_{2}=(0,\sqrt{2\pi}l_{B})$, the wavefunction $|\Phi_{0}\rangle$} with eigenvalue one for both $\hat{T}_{1}$ and $\hat{T}_{2}$ can be written as
@@
due to linearlity (the $\approx$ denotes coarse-graining and ignoring higher-order terms in $\left|\bm{k}\right|^{2}l_{B}^{2}$). This motivates a mathematical map between NC functions and regular functions, known as the Weyl map [DouglasNekrasov2001]. The Weyl map is a linear, invertible map given by
@@ III Many body bosonic coherent states in the LLL
Note that the projection to the LLL comes for free as the coherent states are LLL-restricted, and that the NC field $\hat{f}$ describes a boson. The last step is paramount. The time derivative and chemical potential terms match exactly when we rewrite it in terms of the trace over non-commutative space, but the quartic term cannot be exactly mapped in that way. From the long-wavelength behavior of the wavefunction products (Appendix A and discussions earlier), we know that it is approximated by the leading order term written above, where $g^{\ast}(\approx 1.160g$ for the triangular vortex lattice [YangMacdonald2004], details in Appendix C) is a renormalized interaction vertex, which can be computed by numerically estimating the interaction energy of the reference state $\sim\int d^{2}\bm{r}\left|\Phi_{0}(\bm{r})\right|^{4}$. Since this is not an exact mapping, there are corrections that have to be quartic in the field operators because of the scaling with density, and they come in the form of higher derivative terms in the NC field theory language. These corrections depend on the exact vortex lattice reference state $|\Phi_{0}\rangle$} that we start with, and we will not try to calculate the coefficients from microscopics. Rather, we use symmetry and dimensional analysis to constrain the form of the correction terms.
@@
We find that there is no derivative four-field term that both respect the spatial symmetries and gives an energy cost to unitary $\hat{f}$. There is a similar issue if one tries to write down a local field theory in terms of the boson field operators for a dipole conserving Bose-Hubbard model [Lake2022] which captures the effects of dipole hopping. Another approach is to start from a non-local [^3] Non-local here just means that the terms in the Lagrangian are not simply a function of a field and its derivatives at the same spatial point. The theory is still physically local, it has exponential decay of interactions. hopping term in the Lagrangian
@@
where $g(\bm{a},\bm{b})=g\delta(\bm{a}\cdot\bm{b})\mathcal{O}(e^{-(\bm{a}^{2}+\bm{b}^{2})l_{B}^{-2}})$ is a decaying function reflecting the locality of the hopping process. This form of the hopping term is constrained by magnetic translation and rotation symmetry, which implies dipole moment $\int\rho\hat{r}_{i}$ and trace of quadrupole moment $\int\rho\hat{r}_{i}\hat{r}_{i}/2$ conservation respectively, as these are the generators of those symmetries [Spodyneiko2023]. Then, we use Hubbard-Stratonovich fields (Appendix B) to resolve the interaction terms into quadratic pieces. Because of the symmetries of the problem, we have to introduce two sets of Hubbard-Stratonovich fields, one for the trace of quadrupole moment and a vector field for the dipole moment. Using this, we can systematically form condensates of each of these fields. The mean-field phases naively correspond to
@@
    Crystalline phase: a phase with spontaneously broken rotation and translation symmetry, with a quadratically dispersing NGB. The naive expectation of two linearly dispersing NGB is subverted due to the symmetry allowed $\epsilon_{ij}d_{i}\partial_{\tau}d_{j}$ term, where $e^{id_{j}}$ is the two-component dipole/translation order parameter. This is similar to the NGB physics of a Heisenberg ferromagnet, and is a type-B NGB [WatanabeMurayama2013, Beekman2019].
@@
The naive mean-field theory predicts that there is a cascade of transitions between the trivial insulator and the LLL superfluid via spontaneous breaking of various higher moments of charge conservation symmetry. This naive treatment is wrong, and there is a simple reason why. Take the crystalline phase, for example. This phase can be thought of as a dipole condensate, that is, a phase where dipoles are proliferated starting from the parent trivial insulator. A necessary condition for this to be true is that the average charge density of this state is the same as the trivial insulator phase – which has zero density. But we know that the bosonic vacuum is the only state with exactly zero density, and there are no negative density states [^4] An identical issue arises around the zero filling Mott insulator lobe of the dipolar Bose-Hubbard model and is discussed in footnote 2 of [Lake2022].. Using an analogous argument for the hexatic phase (which is a condensate of quadrupoles), we conclude that there are no non-trivial hexatic and crystalline phases in this problem, and we generically expect a first-order transition between the trivial insulator phase and the LLL superfluid phase. Refs. [NguyenMoroz2023, Du2024b] study these phases from the dual perspective, i. e. in terms of melting of the vortex lattice. Under the identification of the vortex‐liquid with the trivial phase introduced above, their findings appear to contradict our results. Ref. [Du2024b] describes a series of transitions connecting the vortex solid (LLL superfluid) to the vortex liquid (trivial) phase, via the intermediate phases described above. This disagrees with our results as we argue that these intermediate phases do not exist, which becomes apparent when viewed from the original (non-dual) language. Ref. [NguyenMoroz2023] speculates about a continuous direct transition between the vortex solid and vortex liquid phase, but we argue that such a transition will generically be first order. This is because from our perspective, the vortex liquid phase is the same as the bosonic vacuum, and one cannot have a continuous transition between states with different boson density. The disagreement is resolved when we note that the ideas in Refs. [NguyenMoroz2023, Du2024b] become relevant when we observe transitions between bosonic quantum Hall states and LLL superfluids, where the quantum Hall state is a vortex liquid phase at non-zero boson density, distinct from the bosonic vacuum state we consider here. We leave a microscopic understanding to the effective field theory describing this transition to future work.
@@ IV Non-commutative field theory of the LLL superfluid
With the form of the effective non-commutative Lagrangian for the interacting bosons in the LLL, we are ready to reproduce the effective NC field theory action for the LLL superfluid described in Ref. [Du2024a] by integrating out the amplitude mode. Like discussed earlier, we will use some symmetry and dimensional analysis arguments to arrive at the term that provides a dispersion for the Tkachenko mode.
@@
Focusing on the action, we note that the first term is invarant under time-reparametrization symmetry $\hat{u}(\tau)\mapsto\hat{u}(\alpha(\tau))$, where $\alpha$ is a differentiable map. This means that the first term is a geometric phase, since it only depends on the path we take and not how quickly we traverse the path. This is the non-commutative Berry Lagrangian term, discussed previously in a commutative field theoretic description of a LLL superfluid [MorozSon2019]. It encodes the Aharanov-Bohm phase picked up by the superfluid bosons due to the magnetic field. One way to see that is by noting that $\hat{u}$ is a unitary map in the LLL, which in the Heisenberg picture encodes an area preserving deformation (APD) of space [Du2022], given by
@@
That is, we compute the energy cost $\int d^{2}\bm{r}gn_{0}^{2}\left|\Phi_{\hat{u}}(\bm{r})\right|^{4}$ numerically and match it with the effective field theory to get the correct coefficients. The details are shown in Appendix C.2. This quadratic low-energy action is valid as long as the system is short-range interacting, and at the thermodynamic and LLL limit. The Tkachenko field $\hat{\varphi}$ describes both the coarse-grained phase fluctuations of the LLL superfluid, and (its gradient describes) phonon fluctuations of the vortex lattice [Du2024a]. The Berry Lagrangian term is a total derivative term at quadratic order, and does not affect the dispersion relation of the Tkachenko mode, but this term is crucial for the transport physics of the LLL superfluid [Du2024b]. We have reproduced the quadratic EFT described in Ref. Du2024a, along with its coefficients in terms of the microscopic parameters. The dispersion relation
@@
We end this section by noting that the quadratic dispersion of the Tkachenko mode implies that there is no long-range order of $e^{i\hat{\varphi}}$ in the 2+1D LLL superfluid, due to a generalized Hohenberg-Mermin-Wagner theorem [YuanChenYe2020, StahlLakeNandkishore2022, KapustinSpodyneiko2022, Spodyneiko2023]. Instead, there is quasi-long range order and the field theory for the Tkachenko mode describes the spin-wave type excitations similar to the 1+1D superfluid, ignoring topological defects. Although the Tkachenko modes also describe the phonon motion of the vortex lattice (note that translation is spontaneously broken unlike $U(1)$ charge conservation), we believe it is inappropriate to call it a Nambu-Goldstone boson.